OpenStudy (anonymous):

A store allows customers to fill their own bags of candy. Troy decides he only wants chocolate-covered pretzels and gumdrops. Chocolate-covered pretzels sell for $0.89 per pound, and gumdrops sell for $0.65 per pound. Troy’s bag weighs 1.8 pounds and it cost $1.29. Write an equation in simplest form that represents the problem described here. A. 0.24p + 1.17 = 1.8 B. 0.24p + 1.17 = 1.29 C. –0.24p + 1.602 = 1.29 D. –0.24p + 1.602 = 1.8

OpenStudy (ttp):

put amount of chocolate covered pretzels as p, then since the total is 1.8 pound the amount of gumdrops is 1.8- p pretzels cost .89 per pound so p pounds cost .89p and the rest is .65(1.8 - p) they total $1.29 so you know .89p+.65(1.8−p)=1.29 multiply out and combine like terms and you get .24p+1.17=1.29

OpenStudy (anonymous):

thank you

OpenStudy (ttp):

Welcome(:

OpenStudy (anonymous):

A store allows customers to fill their own bags of candy. Terri decides she only wants jelly beans and chocolate drops. Jelly beans sell for $0.98 per pound, and chocolate drops sell for $0.67 per pound. Terri’s bag weighs 2.1 pounds and it costs $1.56. If p = pounds of jelly beans, which equation in simplest form represents the situation described.

OpenStudy (anonymous):

0.31p + 1.41 = 1.56 0.31p + 1.41 = 2.1 –0.31p + 2.06 = 1.56 –0.31p + 2.06 = 2.1 ?

OpenStudy (anonymous):

someone help me please?